# Photometric and spectroscopic variability of blue supergiant $\rho$ Leo

### v1

@@ Abstract
Aims. We investigate the variability patterns of the evolved B-type star $\rho$ Leo to determine its properties, identify the underlying physical processes, and constrain its evolutionary stage. We have combined extensive, long-term datasets of spectroscopic and photometric observations from various sources. These include data from the TESS and Kepler space telescopes, as well as observations from the 1.5 m telescope in Estonia.
Methods. We analysed the data using the Generalized Lomb-Scargle periodogram, Lomb-Scargle periodogram with pre-whitening, and the Weighted Wavelet Z-Transform. To determine the fundamental parameters of $\rho$ Leo, we fitted synthetic line profiles computed with FastWind code to the HARPS spectrum. We used the zpektr code to infer the stellar rotation inclination angle.
@@ 1 Introduction
Over the past few decades, asteroseismology has become a powerful tool for probing stellar interiors. Studying the variability of blue supergiants (BSGs) is particularly important, as their pulsational behaviour provides direct clues about the internal structure and the physical processes that shape the evolution of massive stars (see the review by 2023Ap&SS.368..107B). Investigating objects such as $\rho$ Leo therefore contributes to the broader effort to link observed surface variability with the underlying physics of stellar interiors.
$\rho$ Leo (HD 91316) is a slowly rotating BSG of spectral class B1 Iab (1997yCat..72840265H). It has a mass of approximately 22 $M_{\odot}$, a radius of 32 $R_{\odot}$, an effective temperature of $T_{\rm eff}=20\,260$ K and a projected rotational velocity, $v\sin{i}$=49=49 km s-1(2014A&A...562A.135S). $\rho$ Leo is classified as an $\alpha$ Cygni-type variable, exhibiting quasi-periodic photometric and spectroscopic variability on timescales of days (e.g. 1988ApJS...68..319L, refId1). The binary nature of this star has been extensively discussed in the literature (e.g. 1976A&A....48..245D, 2024AJ....168...28T and references therein); however, no conclusive evidence has been found.
The stellar wind parameters of $\rho$ Leo, including mass loss and wind variability, have been studied by 2006A&A...446..279C. In addition, it is known to display non-radial pulsations (2016MNRAS.458.1604K). The connection between the pulsation pattern and the evolutionary stage of BSGs has been established by 2013MNRAS.433.1246S, yet detailed information on the variability of $\rho$ Leo, the existing periods, and a comprehensive comparison of these periods is still lacking.
Long-term spectroscopic observations of hot stars such as $\rho$ Leo reveal diverse structures within their atmospheres, with lifetimes ranging from fractions of an hour to several days. For OB stars, variations in line profiles are often regular or quasi-regular, with long-term periodic variations spanning several days to tens of days. These periodic changes are frequently associated with variations in mass loss, reflecting the dynamic nature of stellar atmospheres (1980S&T....60R.418C; 1989nos..book.....U; 2010aste.book.....A). Longer periods are more likely to be associated with radial pulsation modes. To investigate this further, 2013MNRAS.433.1246S extended their calculations of pulsation instabilities, modelling pulsation patterns that follow the evolution of massive stars from the main sequence to the red supergiant (RSG) stage and beyond. They conducted these calculations for stars with initial masses up to 25 $M_{\odot}$, considering both non-rotating stars and stars with initial rotation rates of 40% of their critical rotation velocity.
@@
A detailed variability analysis of $\rho$ Leo was carried out by 10.1093/mnras/sty308, who combined 80 days of high-cadence K2 halo photometry with 1800 days of high-resolution HERMES spectroscopy. All detected variability was confined to frequencies below 1.5 d-1. The dominant frequency, $f_{\rm rot}=0.0373$ d-1 (period ~26.8 days), was attributed to rotational modulation at the base of an aspherical stellar wind and interpreted as the stellar rotation signature. In addition, the combined photometric and spectroscopic data revealed low-amplitude photospheric velocity variations in the range ~0.2–0.6 d-1, which were attributed to large-scale, travelling gravity waves in the stellar envelope.
@@
In this study, we used the Weighted Wavelet Z-transform (WWZ) (1996AJ....112.1709F) and Lomb-Scargle analysis methods (1976Ap&SS..39..447L; 1982ApJ...263..835S; 2009A&A...496..577Z) to identify frequencies in the light curves from the Kepler/K2111https://science.nasa.gov/mission/kepler/ and TESS [^2] https://tess.mit.edu/ space telescopes, as well as in our spectroscopic observations. The 2017 spectra were obtained just before the K2 campaign, while in 2022, one spectrum preceded, two coincided with, and the rest closely followed the TESS observations.
In addition to analysing the temporal variability, we employ detailed computational modelling to constrain the inclination of the stellar rotation axis, a parameter that has not been determined for this star in previous studies. Together, these approaches provide new insights into the complex variability of $\rho$ Leo and help to distinguish stable periodic signals from those that evolve with time.
@@ 2.1 TESS photometry
TESS observed $\rho$ Leo from 6 November to 30 December 2021, in sectors 45 and 46, in the short-cadence mode (2 min exposures), for a total of almost 7000 measurements covering 52.44 days. We extracted the light curve of $\rho$ Leo from the TESS full-frame images downloaded from the MAST [^3] https://archive.stsci.edu/missions-and-data/tess archive. Due to the target‘s saturating magnitude, $\rho$ Leo flux was distributed over a relatively large number of TESS pixels. Therefore, the corresponding aperture for the stellar signal and background was carefully selected according to the guidelines provided by 2021AJ....162..170H and 2021ApJS..257...53L. Our custom pixel mask (aperture) for $\rho$ Leo in TESS sector 45 is shown in Fig. 15 in Appendix A. Within this aperture (radius $\sim 2^{\prime}$), there are eight additional faint stars (G = 15.4–20.3 mag, Gaia DR3) that have no measurable effect on the $\rho$ Leo light curve. The stellar signal was measured within the aperture marked by pure-white pixels with counts exceeding approximately four times the background level.
We cannot rule out the possibility that the target is also variable on time scales longer than the duration of a single sector ($\sim 26$ days). Hence, to better match the instrumental time series of both sectors, we also extracted the light curve of the comparison star 48 Leo, which is positioned inside the same CCD field as $\rho$ Leo. Consequently, we present the photometric TESS time series of $\rho$ Leo (Fig. 16) as differential magnitudes against the stable comparison star 48 Leo which has a flux scatter of less than $\sim 1$ ppt.
One of the most striking features of the TESS light curve is its distinct ”sawtooth” shape. Interestingly, the same pattern has been observed in K2 photometry (10.1093/mnras/sty308, Fig. 1), further supporting its validity.
@@ 2.2 Kepler K2 photometry
$\rho$ Leo was observed with the Kepler space telescope in the framework of the K2 Campaign 14 (K2) from 30 May to 19 August 2017, just after our spectroscopic observations in the Season 2017 (Sect. 2.3). The K2 dataset consists of long-cadence data (29.4 min) covering 79.64 d. For our analysis, we used the K2 halo photometric light curve of $\rho$ Leo published by 10.1093/mnras/sty308. The light curve was digitized from their Figure 1, which presents the brightness variations extracted using the first photometric mask (upper left panel in their Figure A1). This is the same aperture configuration that 10.1093/mnras/sty308 employed for their frequency analysis, thereby ensuring that our study is based on an identical dataset and allowing a direct comparison of the results.
@@ 2.3 Spectroscopy
Archival photometric data were complemented with the spectroscopic observations obtained with the Tartu Observatory’s (TO) 1.5 m telescope AZT-12. Long-slit spectrograph ASP-32 in the Cassegrain focus was used with a 1800 lines mm-1 diffraction grating providing spectra in the wavelength range from 6300 Å to 6730 Å with a SNR $\sim 300-400$ and resolution $R\approx 10\,500$. In total, data were collected during 150 observing nights between 2014 and 2024, resulting in approximately 2700 spectra.
A typical sequence of observations covers a duration of 2–4.5 hours during the night with a typical exposure time of approximately 300 seconds (ranging from 150 to 800 seconds depending on weather conditions). In this article, we focus on the two most important observational Seasons: winter/spring 2017 (referred to as Season 2017) and autumn 2021/spring 2022 (referred to as Season 2022). The data for season 2017 were collected just before the K2 photometry, between 4 January and 16 May 2017, comprising a total of 369 spectra over 20 nights. Spectra for Season 2022 were obtained between 1 November 2021 and 15 May 2022, totalling 1106 spectra over 39 nights, collected mostly shortly after the TESS photometry. We note that one spectrum was obtained before the TESS photometry, on 1 November 2021, and two spectra during TESS observations, on 21 November 2021 and 9 December 2021. The data were reduced using the IRAF [^4] IRAF was written at the National Optical Astronomy Observatory, which was operated by the Association of Universities for Research in Astronomy (AURA) under cooperative agreement with the National Science Foundation. [^5] https://iraf-community.github.io/ software (1986SPIE..627..733T; 1993ASPC...52..173T), employing standard routines from the noao, imred and ccdred packages. First, bias and flat-field corrections were applied to the spectra, while dark correction was not necessary because of adequate cooling of the CCD. Next, wavelength calibration was performed using the ThAr lamp frames obtained before and after the target exposures. To further check and refine the stability of the wavelength scale during the observing run, telluric lines near 6570 Å were used. Additionally, heliocentric correction was applied to the spectra. Continuum normalization was achieved by fitting a first-order (cubic spline) function. A summary of the observing periods and the number of spectroscopic data points for all datasets used in this work is given in Table 1, while the full log of observations is presented in Table 8 in Appendix B.
@@
In addition to our own long-term spectroscopic observations, the pipeline-reduced high-resolution spectrum, R = 115 000, obtained with the High Accuracy Radial Velocity Planet Searcher (HARPS) attached to the 3.6 m telescope at La Silla Observatory, observed on 12 February 2006 (HJD-2453778.618849) was downloaded from the ESO Science Archive Facility [^6] Based on data obtained from the ESO Science Archive Facility with DOI https://doi.eso.org/10.18727/archive/33 under the ESO program 60.A-9036. The spectrum has a SNR $\sim 300$ and covers a wavelength range of 378169133781-6913 Å.
@@ 3.1 Stellar parameters of $\rho$ Leo
To reassess the stellar parameters of $\rho$ Leo, the high-resolution HARPS spectrum observed on 12 February 2006 was used. This particular spectrum was selected because the spectral lines of interest appeared symmetric, indicating a period of weak stellar-wind activity. Choosing a spectrum obtained under such calm wind conditions minimises the influence of wind-related distortions on the line profiles and thus ensures the most reliable determination of the atmospheric parameters.
First, we derived the parameters responsible for the broadening of the spectral lines, the macroturbulent velocity, $v_{\rm mac}$ and the projected rotational velocity, $v{\sin i}$. For this purpose, we analysed the silicon line, Si III $\lambda$4552, using the iacob-broad tool [^7] https://research.iac.es/proyecto/iacob/pages/en/useful-tools.php, developed within the framework of the IACOB project (2014A&A...562A.135S). The chosen diagnostic line is well isolated, exhibits a symmetric photospheric profile, and shows minimal wind interference, making it particularly suitable for this analysis.
The iacob-broad tool combines a Fourier transform and a goodness-of-fit analysis to disentangle rotational and macroturbulent broadening, adopting a radial–tangential macroturbulent profile with equal radial and tangential components, following 2014A&A...562A.135S. The code provides two independent estimates of the projected rotational velocity, $v{\sin i}$(FT) and $v{\sin i}$(GOF), whose agreement serves as an internal reliability check. The tool also provides uncertainties from the GOF analysis based on the corresponding $\chi^{2}$ distributions, and these values are reported in Table 2.
Next, we used a set of spectral lines to determine the rest of the stellar parameters with the FastWind v10.6 code (Puls). Uncertainties in $T_{\rm eff}$, $\log g$, $M$ $M_{\odot}$], and $v_{\rm mic}$were estimated by varying the derived parameters with FastWind. The fit to the observed spectrum is shown in Fig. 1.
The results of our computations and measurements, along with previously established parameters, are presented in Table 2.
@@
Figure 1: Comparison between observed spectrum (black line) and the best-fitting FastWind model (red line).
3.2 Time-series analysis
Due to uneven sampling and gaps in the analysed time series, the following widely used techniques were employed in this study to perform the frequency analysis: Generalized Lomb–Scargle (GLS) periodogram, Generalized Lomb-Scargle periodogram with prewhitening (GLSp) and Weighted Wavelet Z-transform (WWZ).
GLS periodogram (2009A&A...496..577Z), an extension of the classical Lomb–Scargle method (1976Ap&SS..39..447L; 1982ApJ...263..835S), accounts for measurement uncertainties and allows for a floating mean. In GLS the underlying model is a full linear least-squares fit of a sinusoid plus offset, which provides an unbiased treatment of the data errors. As shown by 2009A&A...496..577Z, this approach increases the statistical power of the period search and yields more accurate and robust frequency estimates, especially for uneven sampling and noisy time series. For each dataset, the most significant frequencies were identified from the GLS power spectrum and visually inspected to confirm that they are not artifacts of the time sampling. Significance was assessed via the Lomb–Scargle false-alarm probability (FAP), using a 1% global FAP threshold derived from both the analytic Baluev approximation and a bootstrap with 5000 resamples. In the GLS periodograms, peaks above FAP threshold are considered significant.
GLSp. This approach refines the standard Lomb-Scargle analysis by iteratively removing significant frequencies, enabling the identification of weaker signals.
WWZ (1996AJ....112.1709F). In addition to the GLS periodogram, we also applied time-frequency WWZ method to study the temporal evolution of the detected frequencies. Method is a widely used for analysing periodic signals in time series data (e.g. 2023Galax..11...69A; 2024A&A...689A..35C; 2025MNRAS.tmp.1744W). This technique allows the identification of transient or time-dependent periodicities by computing localized power spectra across sliding time windows.

### v2

# Photometric and spectroscopic variability of the blue supergiant $\rho$ Leo

@@ Abstract
Aims. We investigated the variability patterns of the evolved B-type star $\rho$ Leo to determine its properties, identify the underlying physical processes, and constrain its evolutionary stage. We combined extensive long-term datasets of spectroscopic and photometric observations from various sources. These include data from the TESS and Kepler space telescopes, as well as observations from the 1.5 m telescope in Estonia.
Methods. We analysed the data using the generalized Lomb–Scargle periodogram, the Lomb–Scargle periodogram with pre-whitening, and the weighted wavelet Z-transform. To determine the fundamental parameters of $\rho$ Leo, we fitted synthetic line profiles computed with the FastWind code to the HARPS spectrum. We used the zpektr code to infer the stellar rotation inclination angle.
@@ 1 Introduction
Over the past few decades, asteroseismology has become a powerful tool for probing stellar interiors. Studying the variability of blue supergiants (BSGs) is particularly important as their pulsational behaviour provides direct clues about the internal structure and the physical processes that shape the evolution of massive stars (see the review by 2023Ap&SS.368..107B). Investigating objects such as $\rho$ Leo therefore contributes to the broader effort to link observed surface variability with the underlying physics of stellar interiors.
$\rho$ Leo (HD 91316) is a slowly rotating BSG of spectral class B1 Iab (1997yCat..72840265H). It has a mass of approximately 22 $M_{\odot}$, a radius of 32 $R_{\odot}$, an effective temperature of $T_{\rm eff}=20\,260$ K, and a projected rotational velocity of $v\sin{i}$=49=49 km s-1(2014A&A...562A.135S). $\rho$ Leo is classified as an $\alpha$ Cygni-type variable, exhibiting quasi-periodic photometric and spectroscopic variability on timescales of days (e.g. 1988ApJS...68..319L, refId1). The binary nature of this star has been extensively discussed in the literature (e.g. 1976A&A....48..245D, 2024AJ....168...28T and references therein); however, no conclusive evidence has been found.
The stellar wind parameters of $\rho$ Leo, including mass loss and wind variability, have been studied by 2006A&A...446..279C. In addition, it is known to display non-radial pulsations (2016MNRAS.458.1604K). The connection between the pulsation pattern and the evolutionary stage of BSGs has been established by 2013MNRAS.433.1246S, and yet detailed information on the variability of $\rho$ Leo, the existing periods, and a comprehensive comparison of these periods is still lacking.
Long-term spectroscopic observations of hot stars such as $\rho$ Leo reveal diverse structures within their atmospheres, with lifetimes ranging from fractions of an hour to several days. For OB stars, variations in line profiles are often regular or quasi-regular, with long-term periodic variations spanning several days to tens of days. These periodic changes are frequently associated with variations in mass loss, reflecting the dynamic nature of stellar atmospheres (1980S&T....60R.418C; 1989nos..book.....U; 2010aste.book.....A). Longer periods are more likely to be associated with radial pulsation modes. To investigate this further, 2013MNRAS.433.1246S extended their calculations of pulsation instabilities; they modelled pulsation patterns that follow the evolution of massive stars from the main sequence to the red supergiant (RSG) stage and beyond. They conducted these calculations for stars with initial masses up to 25 $M_{\odot}$, considering both non-rotating stars and stars with initial rotation rates of 40% of their critical rotation velocity.
@@
A detailed variability analysis of $\rho$ Leo was carried out by 10.1093/mnras/sty308, who combined 80 days of high-cadence K2 halo photometry with 1800 days of high-resolution HERMES spectroscopy. All detected variability was confined to frequencies below 1.5 d-1. The dominant frequency, $f_{\rm rot}=0.0373$ d-1 (period ~26.8 days), was attributed to rotational modulation at the base of an aspherical stellar wind and interpreted as the stellar rotation signature. In addition, the combined photometric and spectroscopic data revealed low-amplitude photospheric velocity variations in the range ~0.2–0.6 d-1, which were attributed to large-scale travelling gravity waves in the stellar envelope.
@@
For this study we used the weighted wavelet Z-transform (WWZ) (1996AJ....112.1709F) and Lomb–Scargle analysis methods (1976Ap&SS..39..447L; 1982ApJ...263..835S; 2009A&A...496..577Z) to identify frequencies in the light curves from the Kepler/K2111https://science.nasa.gov/mission/kepler/ and TESS [^2] https://tess.mit.edu/ space telescopes, and in our spectroscopic observations. The 2017 spectra were obtained just before the K2 campaign, while in 2022 one spectrum preceded, two coincided with, and the rest closely followed the TESS observations.
In addition to analysing the temporal variability, we employed detailed computational modelling to constrain the inclination of the stellar rotation axis, a parameter that had not been determined for this star in previous studies. Together, these approaches provide new insights into the complex variability of $\rho$ Leo and help to distinguish stable periodic signals from those that evolve with time.
@@ 2.1 TESS photometry
TESS observed $\rho$ Leo from 6 November to 30 December 2021, in sectors 45 and 46, in the short-cadence mode (2 min exposures), for a total of almost 7000 measurements covering 52.44 days. We extracted the light curve of $\rho$ Leo from the TESS full-frame images downloaded from the MAST [^3] https://archive.stsci.edu/missions-and-data/tess archive. Due to the target’s saturating magnitude, $\rho$ Leo flux was distributed over a relatively large number of TESS pixels. Therefore, the corresponding aperture for the stellar signal and background was carefully selected according to the guidelines provided by 2021AJ....162..170H and 2021ApJS..257...53L. Our custom pixel mask (aperture) for $\rho$ Leo in TESS sector 45 is shown in Fig. 15 in Appendix A. Within this aperture (radius $\sim 2^{\prime}$), there are eight additional faint stars (G = 15.4–20.3 mag, Gaia DR3) that have no measurable effect on the $\rho$ Leo light curve. The stellar signal was measured within the aperture marked by pure-white pixels with counts exceeding approximately four times the background level.
We cannot rule out the possibility that the target is also variable on timescales longer than the duration of a single sector ($\sim 26$ days). Hence, to better match the instrumental time series of the two sectors, we also extracted the light curve of the comparison star 48 Leo, which is positioned inside the same CCD field as $\rho$ Leo. Consequently, we present the photometric TESS time series of $\rho$ Leo (Fig. 16) as differential magnitudes against the stable comparison star 48 Leo, which has a flux scatter of less than $\sim 1$ ppt.
One of the most striking features of the TESS light curve is its distinct sawtooth shape. Interestingly, the same pattern has been observed in K2 photometry (10.1093/mnras/sty308, Fig. 1), further supporting its validity.
@@ 2.2 Kepler K2 photometry
$\rho$ Leo was observed with the Kepler space telescope in the framework of the K2 Campaign 14 (K2) from 30 May to 19 August 2017, just after our spectroscopic observations in Season 2017 (Sect. 2.3). The K2 dataset consists of long-cadence data (29.4 min) covering 79.64 d. For our analysis, we used the K2 halo photometric light curve of $\rho$ Leo published by 10.1093/mnras/sty308. The light curve was digitized from their Figure 1, which presents the brightness variations extracted using the first photometric mask (upper left panel in their Figure A1). This is the same aperture configuration that 10.1093/mnras/sty308 employed for their frequency analysis, thereby ensuring that our study is based on an identical dataset and allowing a direct comparison of the results.
@@ 2.3 Spectroscopy
Archival photometric data were complemented with the spectroscopic observations obtained with the Tartu Observatory’s (TO) 1.5 m telescope AZT-12. The long-slit spectrograph ASP-32 in the Cassegrain focus was used with a 1800 line mm-1 diffraction grating providing spectra in the wavelength range from 6300 Å to 6730 Å with a S/N $\sim 300-400$ and resolution $R\approx 10\,500$. In total, data were collected during 150 observing nights between 2014 and 2024, resulting in approximately 2700 spectra.
A typical sequence of observations covers a duration of 2–4.5 hours during the night with a typical exposure time of approximately 300 seconds (ranging from 150 to 800 seconds depending on weather conditions). For this article we focused on the two most important observational seasons: winter–spring 2017 (referred to as Season 2017) and autumn 2021–spring 2022 (referred to as Season 2022). The data for season 2017 were collected just before the K2 photometry, between 4 January and 16 May 2017, comprising a total of 369 spectra over 20 nights. Spectra for Season 2022 were obtained between 1 November 2021 and 15 May 2022, totalling 1106 spectra over 39 nights, collected mostly shortly after the TESS photometry. We note that one spectrum was obtained before the TESS photometry, on 1 November 2021, and two spectra during TESS observations, on 21 November 2021 and 9 December 2021. The data were reduced using the IRAF [^4] IRAF was written at the National Optical Astronomy Observatory, which was operated by the Association of Universities for Research in Astronomy (AURA) under cooperative agreement with the National Science Foundation. [^5] https://iraf-community.github.io/ software (1986SPIE..627..733T; 1993ASPC...52..173T), employing standard routines from the noao, imred, and ccdred packages. First, bias and flat-field corrections were applied to the spectra; dark correction was not necessary because of adequate cooling of the CCD. Next, wavelength calibration was performed using the ThAr lamp frames obtained before and after the target exposures. To further check and refine the stability of the wavelength scale during the observing run, telluric lines near 6570 Å were used. Additionally, heliocentric correction was applied to the spectra. Continuum normalization was achieved by fitting a first-order (cubic spline) function. A summary of the observing periods and the number of spectroscopic data points for all datasets used in this work is given in Table 1, while the full log of observations is presented in Table 8 in Appendix B.
@@
In addition to our own long-term spectroscopic observations, the pipeline-reduced high-resolution spectrum, R = 115 000, obtained with the High Accuracy Radial Velocity Planet Searcher (HARPS) attached to the 3.6 m telescope at La Silla Observatory, observed on 12 February 2006 (HJD-2453778.618849) was downloaded from the ESO Science Archive Facility. [^6] Based on data obtained from the ESO Science Archive Facility with DOI https://doi.eso.org/10.18727/archive/33 under the ESO programme 60.A-9036 The spectrum has a S/N $\sim 300$ and covers a wavelength range of 378169133781-6913 Å.
@@ 3.1 Stellar parameters of $\rho$ Leo
To reassess the stellar parameters of $\rho$ Leo, the high-resolution HARPS spectrum observed on 12 February 2006 was used. This particular spectrum was selected because the spectral lines of interest appeared symmetric, indicating a period of weak stellar-wind activity. Choosing a spectrum obtained under such calm wind conditions minimizes the influence of wind-related distortions on the line profiles, and thus ensures the most reliable determination of the atmospheric parameters.
First, we derived the parameters responsible for the broadening of the spectral lines; the macroturbulent velocity, $v_{\rm mac}$; and the projected rotational velocity, $v{\sin i}$. For this purpose, we analysed the silicon line, Si III $\lambda$4552, using the iacob-broad tool, [^7] https://research.iac.es/proyecto/iacob/pages/en/useful-tools.php developed within the framework of the IACOB project (2014A&A...562A.135S). The chosen diagnostic line is well isolated, exhibits a symmetric photospheric profile, and shows minimal wind interference, making it particularly suitable for this analysis.
The iacob-broad tool combines a Fourier transform and a goodness-of-fit analysis to disentangle rotational and macroturbulent broadening, and adopts a radial–tangential macroturbulent profile with equal radial and tangential components, following 2014A&A...562A.135S. The code provides two independent estimates of the projected rotational velocity, $v{\sin i}$(FT) and $v{\sin i}$(GOF), whose agreement serves as an internal reliability check. The tool also provides uncertainties from the GOF analysis based on the corresponding $\chi^{2}$ distributions, and these values are reported in Table 2.
Next, we used a set of spectral lines to determine the rest of the stellar parameters with the FastWind v10.6 code (Puls). Uncertainties in $T_{\rm eff}$, $\log g$, $M$ $M_{\odot}$], and $v_{\rm mic}$were estimated by varying the derived parameters with FastWind. The fit to the observed spectrum is shown in Fig. 1. The results of our computations and measurements, along with previously established parameters, are presented in Table 2.
@@
Figure 1: Comparison between the observed spectrum (black line) and the best-fitting FastWind model (red line).
3.2 Time series analysis
Due to uneven sampling and gaps in the analysed time series, the following widely used techniques were employed in this study to perform the frequency analysis: generalized Lomb–Scargle (GLS) periodogram, generalized Lomb–Scargle periodogram with pre-whitening (GLSp), and weighted wavelet Z-transform (WWZ).
GLS periodogram (2009A&A...496..577Z). This approach is an extension of the classical Lomb–Scargle method (1976Ap&SS..39..447L; 1982ApJ...263..835S); it accounts for measurement uncertainties and allows for a floating mean. In GLS the underlying model is a full linear least-squares fit of a sinusoid plus offset, which provides an unbiased treatment of the data errors. As shown by 2009A&A...496..577Z, this approach increases the statistical power of the period search and yields more accurate and robust frequency estimates, especially for uneven sampling and noisy time series. For each dataset, the most significant frequencies were identified from the GLS power spectrum and visually inspected to confirm that they are not artefacts of the time sampling. The significance was assessed via the Lomb–Scargle false-alarm probability (FAP), using a 1% global FAP threshold derived from both the analytic Baluev approximation and a bootstrap with 5000 resamples. In the GLS periodograms, peaks above the FAP threshold are considered significant.
GLSp. This approach refines the standard Lomb–Scargle analysis by iteratively removing significant frequencies, thus enabling the identification of weaker signals.
WWZ (1996AJ....112.1709F). In addition to the GLS periodogram, we also applied the time-frequency WWZ method to study the temporal evolution of the detected frequencies. The method is a widely used for analysing periodic signals in time series data (e.g. 2023Galax..11...69A; 2024A&A...689A..35C; 2025MNRAS.tmp.1744W). This technique allows the identification of transient or time-dependent periodicities by computing localized power spectra across sliding time windows.